OpenStudy (anonymous):

Given x/12=5/4, which is proportional to x+5/x-5? (RATIO HELP!)

OpenStudy (anonymous):

\[\frac{x}{12}=\frac{5}{4}\iff x=\frac{12\times 5}{4}=3\times 5=15\] \[\frac{15+5}{15-5}=\frac{20}{10}=2\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw